# Exponentially weighted estimands and the exponential family: filtering, prediction and smoothing

### v1

@@ 7 Conclusions
The exponentially weighted estimands combined with the canonical exponential family delivers simple exact filters, predictors and smoothers. They are in the same statistical spirit as generalized linear models (e.g. McCullaghNelder(89)) and are relatively easy to fit to data using a quasi-likelihood approach either all at once or using a two step approach.
The exponentially weighted estimands framework can be used on non-exponentially family models and even for general loss functions. In such cases the optimization has to be carried out numerically or through a sequential approximation. This then relates to work on stochastic gradient descent (e.g. RobbinsMunro(51), toulis2015scalable; toulis2017asymptotic, toulis2016towards, toulis2021proximal), LangeOsDijk(24), HeelLangeOsDijk(24), as well as the DCS/GAS filter of Harvey(13) and CrealKoopmanLucas(13).
8 Appendix: some proofs
8.1 Proof of Theorem 1
Focus on the filtering case. Now
Differentiate $Q_{\lambda,t}(\theta)$ with respect to $\theta$ and solve. This yields the stated result.
The same argument applied to prediction and smoothing.
The remaining results follow by invariance.

### v2

@@ 7 Conclusions
The exponentially weighted estimands combined with the CEF delivers simple exact filters, predictors and smoothers. They are in the same statistical spirit as generalized linear models (e.g. McCullagh and Nelder (1989)) and are relatively easy to fit to data using a quasi-likelihood approach either all at once or using a two step approach.
The exponentially weighted estimands framework can be used on non-exponentially family models and even for general loss functions. In such cases the optimization has to be carried out numerically or through a sequential approximation.
Acknowledgments
We are grateful for the comments and encouragement from Dick van Dijk, Siem Jan Koopman, Rutger-Jan Lange and Bram van Os. Donker van Heel conducted this research during a visiting research appointment at Harvard University’s Department of Statistics.
Disclosure Statement
Donker van Heel thanks the Fulbright Program for financial support under their scholarship program. The authors report there are no competing interests to declare.